OpenStudy (anonymous):

Simplify the radical. the square root of 90 A. 5 square root of 5 B. 25 square root of 5 C. 10 square root of 5 D. 5 square root of 25

OpenStudy (fretje):

sqrt (90) = 3 * sqrt (10)

OpenStudy (anonymous):

square root of 125* my bad

OpenStudy (anonymous):

5sqrt5

OpenStudy (fretje):

sqrt (125) = sqrt (5²*5) = 5* sqrt (5)

OpenStudy (anonymous):

$$5 \sqrt{5} $$